# Quadratic dispersion relations in gapless frustration-free systems

### v1

@@ III.1 Rank 1 case
where $|\psi\rangle_{x,x+1}$ is a normalized state for two spins at $x$ and $x+1$. Without loss of generality one can assume the form
$|\psi\rangle=(\alpha+i\beta)|01\rangle+(\alpha+i\gamma)|10\rangle+\delta|11\rangle,$ (20)
$\alpha,\beta,\gamma,\delta\in\mathbb{R},\quad 2\alpha^{2}+\beta^{2}+\gamma^{2}+\delta^{2}=1$ (21)
by performing a local unitary transformation (see Appendix B for the derivation) Bravyi and Gosset (2015). The model has U(1) spin rotation symmetry about the $z$ axis if and only if $\delta=0$.
Since the matrix elements are known, the Hamiltonian in Eq. (19) can be expressed in terms of spin operators $\hat{s}_{x}^{a}$ and $\hat{s}_{x+1}^{a}$ ($a=x,y,z$). For example, when $(\alpha,\beta,\gamma,\delta)=\pm(0,1/\sqrt{2},-1/\sqrt{2},0)$, $\hat{Q}_{x,x+1}$ becomes the projector onto the spin-singlet state $(|01\rangle-|10\rangle)/\sqrt{2}$:
$\hat{Q}_{x,x+1}$ $=\hat{\Pi}_{x,x+1}^{s=0},$ (22)
$\hat{\Pi}_{x,x+1}^{s=0}$ $\coloneqq\hat{\mathbbm{1}}-\frac{1}{2}(\hat{\bm{s}}_{x}+\hat{\bm{s}}_{x+1})^{2}=\frac{1}{4}\hat{\mathbbm{1}}-\hat{\bm{s}}_{x}\cdot\hat{\bm{s}}_{x+1}.$ (23)
The corresponding Hamiltonian is the ferromagnetic Heisenberg model. Any state that is invariant under permutations of any two qubits (e.g., the fully-polarized state and the W state) is a ground state of this model. As a basis of such states, we can choose
$\hat{\mathcal{S}}|\underbrace{0\cdots 0}_{L-n}\underbrace{1\cdots 1}_{n}\rangle\quad(n=0,1,\cdots L),$ (24)
where $\hat{\mathcal{S}}$ is the symmetrization operator, uniformly averaging over the $\binom{L}{n}=\frac{L!}{(L-n)!n!}$ states. This model is known to be gapless due to the presence of a spin-wave excitation $E(k)=2\sin^{2}(k/2)$ under PBC.
To analyze more general cases, let us introduce a matrix
$m_{\psi}\coloneqq\begin{pmatrix}\langle\psi|01\rangle&\langle\psi|11\rangle\\ -\langle\psi|00\rangle&-\langle\psi|10\rangle\end{pmatrix}=\begin{pmatrix}\alpha-i\beta&\delta\\ 0&-\alpha+i\gamma\end{pmatrix}.$ (25)
The model defined by Eqs. (14), (19), and (20) can be gapless only when $\det m_{\psi}=(\alpha-i\beta)(-\alpha+i\gamma)\neq 0$, i.e., $|\psi\rangle$ is entangled Bravyi and Gosset (2015). In this case, we introduce an invertible operator $\hat{M}$ composed of nonuniform powers of the matrix $m_{\psi}$:
@@
This operator maps $|\psi\rangle$ to the singlet state:
$\hat{M}^{\dagger}|\psi\rangle_{x,x+1}=(\det{m_{\psi}}^{*})^{j}(|01\rangle_{x,x+1}-|10\rangle_{x,x+1}),$ (28)
It follows that $\hat{Q}_{x,x+1}$ for the state in Eq. (20) can be connected to $\hat{\Pi}_{x,x+1}^{s=0}$ as
@@
This type of non-unitary transformation is known as Witten’s conjugation Witten (1982); Wouters et al. (2021). As a consequence, ground states of $\hat{H}^{\mathrm{obc}}$ takes the form Bravyi and Gosset (2015)
$\hat{M}\hat{\mathcal{S}}|\underbrace{0\cdots 0}_{L-n}\underbrace{1\cdots 1}_{n}\rangle\quad(n=0,1,\cdots L).$ (30)
The Hamiltonian under PBC has an additional term $\hat{Q}_{L,1}$. Every ground state of $\hat{H}^{\mathrm{obc}}$ remains a ground state of $\hat{H}^{\mathrm{pbc}}$ if $m_{\psi}^{L}$ is proportional to the identity matrix $\mathbbm{1}$. However, except when $(\alpha,\beta,\gamma,\delta)=\frac{\pm 1}{\sqrt{2}}(0,1,-1,0)$ discussed above, $m_{\psi}^{L}$ is not proportional to $\mathbbm{1}$ for some $L$. In such a case, the ground state of $\hat{H}^{\mathrm{pbc}}$ is restricted to the product state of the eigenstates of $\hat{m}_{\psi}$; that is, $|{\Phi_{0}}\rangle=|00\cdots 0\rangle$ and $|uu\cdots u\rangle$, where
$|u\rangle\coloneqq\frac{[2\alpha-i(\beta+\gamma)]|1\rangle-\delta|0\rangle}{\sqrt{4\alpha^{2}+(\beta+\gamma)^{2}+\delta^{2}}}.$ (31)
One can easily check that $\langle\psi|00\rangle=\langle\psi|uu\rangle=0$. Therefore, the exact ground state degeneracy of this model can be summarized as Bravyi and Gosset (2015)
$\tilde{N}_{\mathrm{deg}}=\begin{cases}L+1&\text{if }m_{\psi}^{L}\propto\mathbbm{1},\\ 2&\text{otherwise}.\\ \end{cases}$ (32)
The variational state $|{\Psi_{k}}\rangle$ in Eq. (16) with $k\neq 0$ is orthogonal to the ground states at least when $m_{\psi}^{L}$ is not proportional to $\mathbbm{1}$, since their translation eigenvalues are different. The energy expectation value in Eq. (18) can be computed as
$E(k)$ $=\sum_{x=1}^{L}\big{|}\langle\psi|_{x,x+1}|{\Psi_{k}}\rangle\big{|}^{2}=\big{|}\langle\psi|10\rangle+e^{ik}\langle\psi|01\rangle\big{|}^{2}$
$=\big{|}(\alpha-i\gamma)+(\alpha-i\beta)e^{ik}\big{|}^{2}$
$=(1-\delta^{2})+2(\alpha^{2}+\beta\gamma)\cos k+2\alpha(\beta-\gamma)\sin k$
$=(1-\delta^{2})-\sqrt{(1-\delta^{2})^{2}-(\beta^{2}-\gamma^{2})^{2}}\cos(k-k_{0}).$ (33)
In going to the last line, we defined
$\cos k_{0}\coloneqq-\frac{\alpha^{2}+\gamma\beta}{\sqrt{(\alpha^{2}+\gamma\beta)^{2}+\alpha^{2}(\beta-\gamma)^{2}}},$ (34)
$\sin k_{0}\coloneqq-\frac{\alpha(\beta-\gamma)}{\sqrt{(\alpha^{2}+\gamma\beta)^{2}+\alpha^{2}(\beta-\gamma)^{2}}}.$ (35)
It follows that $E(k)$ can be bounded below as
$E(k)\geq(1-\delta^{2})-\sqrt{(1-\delta^{2})^{2}-(\beta^{2}-\gamma^{2})^{2}}.$ (36)
The right-hand side vanishes when $\beta^{2}=\gamma^{2}$, implying the existence of gapless excitations when $\gamma=\pm\beta$. In this case $E(k)$ can be rewritten as
$E(k)=2(1-\delta^{2})\sin^{2}((k-k_{0})/2),$ (37)
which confirms our conjecture in the rank 1 case. When $\gamma=\beta$, $k_{0}$ is precisely $\pi$, but it takes more general value when $\gamma=-\beta$. This result is consistent with Ref. Bravyi and Gosset (2015) which mathematically proved that $\hat{H}^{\mathrm{obc}}$ is gapped if and only if $\beta^{2}\neq\gamma^{2}$.
@@ III.2 Rank 2 case
Next let us discuss the rank 2 case. In this case the Hamiltonian in Eq. (14) can be gapless only when $\hat{Q}_{x,x+1}$ is unitary equivalent to the XY model with an external magnetic field:
$\hat{Q}_{x,x+1}\coloneqq\frac{1}{2}\hat{\mathbbm{1}}-\frac{\zeta\hat{s}_{x}^{+}\hat{s}_{x+1}^{-}+\zeta^{*}\hat{s}_{x}^{-}\hat{s}_{x+1}^{+}}{1+|\zeta|^{2}}-\frac{|\zeta|^{2}\hat{s}_{x}^{z}+\hat{s}_{x+1}^{z}}{{1+|\zeta|^{2}}}.$ (38)
with $\zeta\in\mathbb{C}$ Bravyi and Gosset (2015). One may express this Hamiltonian as
$\mathrm{ker}\,\hat{Q}_{x,x+1}=\mathrm{span}\big{\{}|00\rangle,|10\rangle+\zeta|01\rangle\big{\}}.$ (39)
@@
When $\zeta=1$, ground states of $\hat{Q}_{x,x+1}^{\zeta=1}$ are the fully-polarized state $|{\Phi_{0}}\rangle$ and the zero momentum state of single spin flip $|\Psi_{0}\rangle$, regardless of the boundary condition. We find that $\hat{Q}_{x,x+1}$ with $\zeta\neq 1$ can be obtained from $\hat{Q}_{x,x+1}^{\zeta=1}$ by Witten’s conjugation as
$\hat{Q}_{x,x+1}=(\hat{M}\hat{Q}_{x,x+1}^{\zeta=1}\hat{M}^{-1})^{\dagger}\hat{C}_{x,x+1}(\hat{M}\hat{Q}_{x,x+1}^{\zeta=1}\hat{M}^{-1}),$ (40)
where
$\hat{M}\coloneqq\hat{\mathbbm{1}}\otimes\hat{m}_{\zeta}\otimes\hat{m}_{\zeta}^{2}\otimes\cdots\otimes\hat{m}_{\zeta}^{L-1},$ (41)
$\hat{m}_{\zeta}\coloneqq|0\rangle\langle 0|+\zeta|1\rangle\langle 1|,$ (42)
$\hat{C}\coloneqq\hat{\mathbbm{1}}+\frac{(1-|\zeta|^{2})^{2}}{2\mathrm{Re}\zeta(1+|\zeta|^{2})}(|10\rangle\langle 01|+|01\rangle\langle 10|).$ (43)
@@
Hence, the ground states of $\hat{H}^{\mathrm{obc}}$ with $\zeta\neq 1$ are given by $\hat{M}|{\Phi_{0}}\rangle=|{\Phi_{0}}\rangle$ and $\hat{M}|\Psi_{0}\rangle\propto\sum_{x=1}^{L}\zeta^{x-1}|x\rangle.$ The latter remains a ground state of $\hat{H}^{\mathrm{pbc}}$ if and only if $\zeta^{L}=1$.
@@
Because of the U(1) symmetry of the Hamiltonian in Eq. (38), the plane-wave state $|{\Psi_{k}}\rangle$ in Eq. (16) is actually an exact eigenstate of $\hat{H}^{\mathrm{pbc}}$. The eigen energy is given by
$E(k)$ $=\frac{|e^{ik}-\zeta|^{2}}{1+|\zeta|^{2}}=\frac{\big{|}e^{i(k-k_{0})}-|\zeta|\big{|}^{2}}{1+|\zeta|^{2}}$
$=\frac{1-|\zeta|^{2}}{1+|\zeta|^{2}}+\frac{4|\zeta|}{1+|\zeta|^{2}}\sin^{2}((k-k_{0})/2),$ (44)
@@
where $k_{0}$ is defined by $\zeta=|\zeta|e^{ik_{0}}$. Hence, the model is gapless and the conjecture holds when $|\zeta|=1$, regardless of $\zeta^{L}=1$ or not.
III.3 Degeneracy in gapped cases
Although our main focus is on gapless models, let us discuss the ground state degeneracy of gapped cases for the model in Eq. (14). This result will be used in our discussion for higher dimensional models in Sec. IV.
Let us start with the rank 1 model. In the $\gamma^{2}\neq\beta^{2}$ case, the formula Eq. (32) implies that $N_{\mathrm{deg}}=2$, because $m_{\psi}^{L}$ cannot be proportional to $\mathbbm{1}$ when the absolute values of the diagonal elements $\alpha-i\beta$ and $-\alpha+i\gamma$ do not agree. The $\alpha=\beta=\gamma=0$ case is exceptional. In this case $|\psi\rangle=\pm|11\rangle$ is a symmetric product state and any product state with no consecutive $1$’s is a ground state. Hence, the ground state degeneracy under PBC with $L\geq 3$ is given by the Lucas number Wikipedia article
$N_{\mathrm{deg}}=\sum_{n=0}^{\lfloor\frac{L+1}{2}\rfloor}\binom{L-n+1}{n}-\sum_{n=2}^{\lfloor\frac{L+1}{2}\rfloor}\binom{L-n-1}{n-2},$ (45)
which increases exponentially with the system size $L$. Here $\lfloor x\rfloor$ represents the greatest integer less than or equal to $x\in\mathbb{R}$. We append the derivation in Appendix C.
Next, let us discuss the rank 2 case. When $|\zeta|\neq 1$, the ground state of $\hat{H}^{\mathrm{pbc}}$ composed of the projector in Eq. (38) is unique ($N_{\mathrm{deg}}=1$) because $\zeta^{L}$ cannot be 1. According to Ref. Bravyi and Gosset (2015), $N_{\mathrm{deg}}\leq 2$ holds even for other rank 2 models under OBC. The min-max principle in Eq. (12) implies the same under PBC.
In summary, for translation-invariant $s=1/2$ chains with nearest neighbor interactions, the ground state degeneracy in the gapped cases is bounded as
$N_{\mathrm{deg}}\leq 2,$ (46)
except for the rank 1 model with $\alpha=\beta=\gamma=0$.
Figure 1: The illustration of the 2D model in Eq. (47) and its decomposition into 1D models.
@@ III.3 Degeneracy in gapped cases
In this section, we generalize the results in the previous section to two and three dimensional models using the min-max principle.
We consider a square lattice of $s=1/2$ spins and assume the Hamiltonian of the form:
$\hat{H}^{\mathrm{2D}}\coloneqq\sum_{x,y=1}^{L}\big{(}\hat{Q}_{(x,y),(x+1,y)}+\hat{Q}_{(x,y),(x,y+1)}\big{)},$ (47)
@@
where $\hat{Q}_{(x,y),(x^{\prime},y^{\prime})}$ is a projector acting on two spins at $(x,y)$ and $(x^{\prime},y^{\prime})$ [Fig. 1 (a)]. We impose PBC and assume the translation invariance in both $x$ and $y$ directions. Our goal is to show that when $\hat{H}^{\mathrm{2D}}$ is frustration free and gapless, there exists a variational state with properties in Eqs. (8) and (9).
Let us decompose the 2D Hamiltonian into 1D chains in two different ways.
$\hat{H}^{\mathrm{2D}}=\hat{H}^{(1)}+\hat{H}^{(2)}=\hat{H}^{\prime}{}^{(1)}+\hat{H}^{\prime}{}^{(2)}.$ (48)
The first choice is the simple one, decomposing the system into decoupled chains as illustrated in Fig. 1 (b) and (c):
$\hat{H}^{(1)}\coloneqq\sum_{y=1}^{L}\hat{H}_{y}^{\mathrm{1D}},\quad\hat{H}_{y}^{\mathrm{1D}}\coloneqq\sum_{x=1}^{L}\hat{Q}_{(x,y),(x+1,y)},$ (49)
$\hat{H}^{(2)}\coloneqq\sum_{x=1}^{L}\hat{H}_{x}^{\mathrm{1D}},\quad\hat{H}_{x}^{\mathrm{1D}}\coloneqq\sum_{y=1}^{L}\hat{Q}_{(x,y),(x,y+1)}.$ (50)
For each $y$, $\hat{H}_{y}^{\mathrm{1D}}$ describes the Hamiltonian for the chain along the $x$ axis; similarly, for each $x$, $\hat{H}_{x}^{\mathrm{1D}}$ describes the Hamiltonian for the chain along the $y$ axis. The second choice is illustrated in Fig. 1 (d) and (e), in which the chains are connected into one piece.
@@
Even when these 1D Hamiltonians are frustration-free, whether the total Hamiltonian $\hat{H}^{\mathrm{2D}}=\hat{H}^{(1)}+\hat{H}^{(2)}=\hat{H}^{\prime}{}^{(1)}+\hat{H}^{\prime}{}^{(2)}$ remains frustration-free or not is, in general, nontrivial. However, this issue can be easily gone around in our case. When either $\hat{H}^{\prime}{}^{(1)}$ or $\hat{H}^{\prime}{}^{(2)}$ is gapped with a finite ground state degeneracy, the min-max principle suggests that $\hat{H}^{\mathrm{2D}}$ is also gapped, because $\hat{H}^{\prime}{}^{(2)}=\hat{H}^{\mathrm{2D}}-\hat{H}^{\prime}{}^{(1)}$ and $\hat{H}^{\prime}{}^{(1)}=\hat{H}^{\mathrm{2D}}-\hat{H}^{\prime}{}^{(2)}$ are positive semi-definite  [^1] This argument cannot be directly applied to the simpler decomposition $\hat{H}^{(1)}+\hat{H}^{(2)}$, because if $N_{\mathrm{deg}}\geq 2$ for each $\hat{H}_{y}^{\mathrm{1D}}$, then the ground state degeneracy of $\hat{H}^{(1)}$ becomes $N_{\mathrm{deg}}^{L}$, which grows exponentially with the system size and the min-max principle becomes silent. For the same reason, the case of $\alpha=\beta=\gamma=0$ in the rank 1 model should be treated separately, but this case is trivially gapped in any spatial dimension. . Hence, we can restrict ourselves to the case where both $\hat{H}^{\prime}{}^{(1)}$ and $\hat{H}^{\prime}{}^{(2)}$ are gapless, implying that $\hat{H}^{(1)}$ and $\hat{H}^{(2)}$ are also gapless. Furthermore, as we have seen in Sec. III, when the Hamiltonian for an $s=1/2$ spin chain with nearest-neighbor interaction is gapless, the fully polarized state $|{\Phi_{0}}\rangle$ in Eq. (15) is a common zero-energy ground state among all models. Hence, the 2D version of the fully polarized state
$|{\Phi_{0}}\rangle\coloneqq\bigotimes_{x,y=1}^{L}|0\rangle_{(x,y)}$ (51)
is a simultaneous ground state of all terms in $\hat{H}^{\mathrm{2D}}$, implying that $\hat{H}^{\mathrm{2D}}$ is frustration-free.
@@
To construct a variational state for low-energy excitations, let us define the plane-wave state of a single spin flip by
$|\Phi_{(k_{x},k_{y})}\rangle\coloneqq\frac{1}{L}\sum_{x,y=1}^{L}e^{i(k_{x}x+k_{y}y)}\hat{s}_{(x,y)}^{-}|{\Phi_{0}}\rangle.$ (52)
The variational energy is given by
$E(k_{x},k_{y})$ $\coloneqq\langle\Phi_{(k_{x},k_{y})}|\hat{H}^{\mathrm{2D}}|\Phi_{(k_{x},k_{y})}\rangle$
$=\sum_{y=1}^{L}\langle\Phi_{(k_{x},k_{y})}|\hat{H}_{y}^{\mathrm{1D}}|\Phi_{(k_{x},k_{y})}\rangle$
$\quad+\sum_{x=1}^{L}\langle\Phi_{(k_{x},k_{y})}|\hat{H}_{x}^{\mathrm{1D}}|\Phi_{(k_{x},k_{y})}\rangle$
$=E(k_{x})+E(k_{y}),$ (53)
where we used the relation
$E(k_{x})=\frac{1}{L}\sum_{x,x^{\prime}=1}^{L}e^{ik_{x}(x-x^{\prime})}\langle x^{\prime},y|\hat{H}_{y}^{\mathrm{1D}}|x,y\rangle,$ (54)
$E(k_{y})=\frac{1}{L}\sum_{y,y^{\prime}=1}^{L}e^{ik_{y}(y-y^{\prime})}\langle x,y^{\prime}|\hat{H}_{x}^{\mathrm{1D}}|x,y\rangle.$ (55)
Therefore,
$E(k_{x},k_{y})=O\big{(}(k_{x}-k_{0})^{2},(k_{y}-k_{0})^{2}\big{)}$ (56)

### v2

@@ III.1 Rank 1 case
where $|\psi\rangle_{x,x+1}$ is a normalized state for two spins at $x$ and $x+1$. Without loss of generality one can assume Bravyi and Gosset (2015)
$|\psi\rangle=(\alpha+i\beta)|01\rangle+(\alpha+i\gamma)|10\rangle+\delta|11\rangle$ (8)
with $\alpha,\beta,\gamma,\delta\in\mathbb{R}$ and $12\alpha^{2}+\beta^{2}+\gamma^{2}+\delta^{2}=$ 1, by performing a local unitary transformation (see Appendix B). For example, when $(\alpha,\beta,\gamma,\delta)=\frac{\pm 1}{\sqrt{2}}(0,1,-1,0)$, $\hat{Q}_{x,x+1}=\hat{\Pi}_{x,x+1}^{s=0}\coloneqq\frac{1}{4}\hat{\mathbbm{1}}-\hat{\bm{s}}_{x}\cdot\hat{\bm{s}}_{x+1}$ is the projector onto the spin-singlet state $|s\rangle\coloneqq(|01\rangle-|10\rangle)/\sqrt{2}$ and the corresponding Hamiltonian is the ferromagnetic Heisenberg model. Any state that is invariant under permutations of any two qubits (e.g., the fully-polarized state and the W state) is a ground state of this model. As a basis of such states, we can choose $\{\hat{\mathcal{S}}|n\rangle\}_{n=0}^{L}$, where $|n\rangle\coloneqq|\underbrace{0\cdots 0}_{L-n}\underbrace{1\cdots 1}_{n}\rangle$ and $\hat{\mathcal{S}}$ is the symmetrization operator, uniformly averaging over the $\binom{L}{n}=\frac{L!}{(L-n)!n!}$ states.
To analyze more general cases, let us introduce a matrix $m_{\psi}\coloneqq\begin{pmatrix}\alpha-i\beta&\delta\\ 0&-\alpha+i\gamma\end{pmatrix}$. The model defined by Eqs. (7), (8) can be gapless only when $\det m_{\psi}=(\alpha-i\beta)(-\alpha+i\gamma)\neq 0$, i.e., $|\psi\rangle$ is entangled Bravyi and Gosset (2015). In this case, we introduce an invertible operator $\hat{M}$ composed of nonuniform powers of the matrix $m_{\psi}$:
@@
This operator maps $|\psi\rangle$ to the singlet state $\hat{M}^{\dagger}|\psi\rangle=(\det{m_{\psi}}^{*})^{j}|s\rangle$. It follows that $\hat{Q}_{x,x+1}$ for the state in Eq. (8) can be connected to $\hat{\Pi}_{x,x+1}^{s=0}$ as
@@
This type of non-unitary transformation is known as Witten’s conjugation Witten (1982); Wouters et al. (2021). As a consequence, ground states of $\hat{H}^{\mathrm{obc}}$ takes the form $\hat{M}\hat{\mathcal{S}}|n\rangle$ ($n=0,1,\cdots L$Bravyi and Gosset (2015).
The Hamiltonian under PBC has an additional term $\hat{Q}_{L,1}$. Every ground state of $\hat{H}^{\mathrm{obc}}$ remains a ground state of $\hat{H}$ if $m_{\psi}^{L}$ is proportional to the identity matrix $\mathbbm{1}$. However, except when $(\alpha,\beta,\gamma,\delta)=\frac{\pm 1}{\sqrt{2}}(0,1,-1,0)$ discussed above, $m_{\psi}^{L}$ is not proportional to $\mathbbm{1}$ for some $L$. In such a case, the ground state of $\hat{H}$ is restricted to the product state of the eigenstates of $\hat{m}_{\psi}$; that is, $|{\Phi_{0}}\rangle=|00\cdots 0\rangle$ and $|uu\cdots u\rangle$, where $|u\rangle\coloneqq\frac{[2\alpha-i(\beta+\gamma)]|1\rangle-\delta|0\rangle}{\sqrt{4\alpha^{2}+(\beta+\gamma)^{2}+\delta^{2}}}$.
The variational state $|{\Psi_{k}}\rangle$ in Eq. (6) with $k\neq 0$ is orthogonal to the ground states at least when $m_{\psi}^{L}$ is not proportional to $\mathbbm{1}$, since their translation eigenvalues are different. The energy expectation value $\langle\Psi_{k}|\hat{H}|\Psi_{k}\rangle$ is given by
$E(k)$ $=\big{|}(\alpha-i\gamma)+(\alpha-i\beta)e^{ik}\big{|}^{2}$
$=(1-\delta^{2})-A\cos(k-k_{0}),$ (12)
where $A\coloneqq\sqrt{(1-\delta^{2})^{2}-(\beta^{2}-\gamma^{2})^{2}}$ and $k_{0}$ is defined by $\cos k_{0}=-2(\alpha^{2}+\gamma\beta)/A$ and $\sin k_{0}=-2\alpha(\beta-\gamma)/A$. This excitation energy takes the minimum value at $k=k_{0}$, which vanishes if $\beta^{2}=\gamma^{2}$, implying the existence of gapless excitations. In this case $E(k)$ can be rewritten as $E(k)=2(1-\delta^{2})\sin^{2}((k-k_{0})/2)$, confirming our conjecture in this model. When $\gamma=\beta$, $k_{0}$ is precisely $\pi$, but it takes more general value when $\gamma=-\beta$.
@@ III.2 Rank 2 case
Next we discuss the rank 2 case. This time the Hamiltonian can be gapless only when $\hat{Q}_{x,x+1}$ is unitary equivalent to the XY model with an external magnetic field Bravyi and Gosset (2015):
$\hat{Q}_{x,x+1}\coloneqq\frac{1}{2}\hat{\mathbbm{1}}-\frac{\zeta\hat{s}_{x}^{+}\hat{s}_{x+1}^{-}+\text{h.c.}}{1+|\zeta|^{2}}-\frac{|\zeta|^{2}\hat{s}_{x}^{z}+\hat{s}_{x+1}^{z}}{{1+|\zeta|^{2}}}$ (13)
@@
with $\zeta\in\mathbb{C}$. This Hamiltonian can also be expressed as $\mathrm{ker}\,\hat{Q}_{x,x+1}=\mathrm{span}\big{\{}|00\rangle,|10\rangle+\zeta|01\rangle\big{\}}$. When $\zeta=1$, ground states of $\hat{Q}_{x,x+1}^{\zeta=1}$ are the fully-polarized state $|{\Phi_{0}}\rangle$ and the zero momentum state of single spin flip $|\Psi_{0}\rangle$, regardless of the boundary condition. We find that $\hat{Q}_{x,x+1}$ with $\zeta\neq 1$ can be obtained from $\hat{Q}_{x,x+1}^{\zeta=1}$ by Witten’s conjugation as
$\hat{Q}_{x,x+1}=(\hat{M}\hat{Q}_{x,x+1}^{\zeta=1}\hat{M}^{-1})^{\dagger}\hat{C}(\hat{M}\hat{Q}_{x,x+1}^{\zeta=1}\hat{M}^{-1}),$ (14)
@@
where $\hat{M}\coloneqq\hat{\mathbbm{1}}\otimes\hat{m}_{\zeta}\otimes\cdots\otimes\hat{m}_{\zeta}^{L-1}$, $\hat{m}_{\zeta}\coloneqq|0\rangle\langle 0|+\zeta|1\rangle\langle 1|$, and $\hat{C}\coloneqq\hat{\mathbbm{1}}+\frac{(1-|\zeta|^{2})^{2}}{2\mathrm{Re}\zeta(1+|\zeta|^{2})}(|10\rangle\langle 01|+|01\rangle\langle 10|)$. Hence, the ground states of $\hat{H}^{\mathrm{obc}}$ with $\zeta\neq 1$ are given by $\hat{M}|{\Phi_{0}}\rangle=|{\Phi_{0}}\rangle$ and $\hat{M}|\Psi_{0}\rangle\propto\sum_{x=1}^{L}\zeta^{x-1}|x\rangle.$ The latter remains a ground state under PBC if and only if $\zeta^{L}=1$.
@@
Because of the U(1) symmetry of the Hamiltonian in Eq. (13), the plane-wave state $|{\Psi_{k}}\rangle$ in Eq. (6) is actually an exact eigenstate of $\hat{H}$. The eigenenergy is
$E(k)=\frac{|e^{ik}-\zeta|^{2}}{1+|\zeta|^{2}}=\frac{1-|\zeta|^{2}}{1+|\zeta|^{2}}+\frac{4|\zeta|}{1+|\zeta|^{2}}\sin^{2}\Big{(}\frac{k-k_{0}}{2}\Big{)},$ (15)
@@
where $k_{0}$ is defined by $\zeta=|\zeta|e^{ik_{0}}$. Hence, the model is gapless and the conjecture holds when $|\zeta|=1$.
Finally, for the gapped cases, we found that the ground state degeneracy is, in general, bounded as $N_{\mathrm{deg}}\leq 2$, except the rank 1 model with $\alpha=\beta=\gamma=0$ for which
$N_{\mathrm{deg}}=\sum_{n=0}^{\lfloor\frac{L+1}{2}\rfloor}\binom{L-n+1}{n}-\sum_{n=2}^{\lfloor\frac{L+1}{2}\rfloor}\binom{L-n-1}{n-2}.$ (16)
See Appendix C for the derivation. This coincides with the Lucas number Wikipedia article that increases exponentially with the system size $L$.
@@
In this section, we generalize results obtained in Sec. III to higher dimensions. We consider a square lattice of $s=1/2$ spins, assuming the Hamiltonian of the form $\hat{H}\coloneqq\sum_{x,y=1}^{L}\big{(}\hat{Q}_{(x,y),(x+1,y)}+\hat{Q}_{(x,y),(x,y+1)}\big{)}$, where $\hat{Q}_{(x,y),(x^{\prime},y^{\prime})}$ is a projector acting on two spins at $(x,y)$ and $(x^{\prime},y^{\prime})$ [Fig. 1 (a)]. We impose PBC and assume the translation invariance in both $x$ and $y$ directions. Our goal is to show that, if $\hat{H}$ is FF and gapless, there exists a variational state $|\Psi_{\bm{k}}\rangle$ with the dispersion satisfying (4).
@@
To this end, let us decompose the 2D Hamiltonian into 1D chains in two different ways. The first choice is the simple one, decomposing $\hat{H}$ into decoupled chains $\hat{H}^{(1)}\coloneqq\sum_{y=1}^{L}\hat{H}_{y}^{\mathrm{1D}}$ and $\hat{H}^{(2)}\coloneqq\sum_{x=1}^{L}\hat{H}_{x}^{\mathrm{1D}}$ as illustrated in Fig. 1 (b,c). For each $y$, $\hat{H}_{y}^{\mathrm{1D}}\coloneqq\sum_{x=1}^{L}\hat{Q}_{(x,y),(x+1,y)}$ describes the Hamiltonian for the chain along the $x$ axis; similarly, for each $x$, $\hat{H}_{x}^{\mathrm{1D}}\coloneqq\sum_{y=1}^{L}\hat{Q}_{(x,y),(x,y+1)}$ describes the Hamiltonian for the chain along the $y$ axis. The second decomposition $\hat{H}=\hat{H}^{\prime}{}^{(1)}+\hat{H}^{\prime}{}^{(2)}$ is illustrated in Fig. 1 (d,e), in which the chains are connected into one piece. Even when these 1D Hamiltonians are FF, whether the total Hamiltonian $\hat{H}$ remains FF or not is, in general, nontrivial. However, this issue can be easily gone around in our case. When either $\hat{H}^{\prime}{}^{(1)}$ or $\hat{H}^{\prime}{}^{(2)}$ is gapped with a finite ground state degeneracy, the min-max principle suggests that $\hat{H}$ is also gapped, because $\hat{H}^{\prime}{}^{(2)}=\hat{H}-\hat{H}^{\prime}{}^{(1)}$ and $\hat{H}^{\prime}{}^{(1)}=\hat{H}-\hat{H}^{\prime}{}^{(2)}$ are positive semi-definite  [^1] This argument cannot be directly applied to the simpler decomposition $\hat{H}^{(1)}+\hat{H}^{(2)}$, because if $N_{\mathrm{deg}}\geq 2$ for each $\hat{H}_{y}^{\mathrm{1D}}$, then the ground state degeneracy of $\hat{H}^{(1)}$ becomes $N_{\mathrm{deg}}^{L}$, which grows exponentially with the system size and the min-max principle becomes silent. For the same reason, the case of $\alpha=\beta=\gamma=0$ in the rank 1 model should be treated separately, but this case is trivially gapped in any spatial dimension. . Hence, we can restrict ourselves to the case where both $\hat{H}^{\prime}{}^{(1)}$ and $\hat{H}^{\prime}{}^{(2)}$ are gapless, implying that $\hat{H}^{(1)}$ and $\hat{H}^{(2)}$ are also gapless. Furthermore, as we have seen above, when the Hamiltonian for an $s=1/2$ spin chain with nearest-neighbor interaction is gapless, the fully polarized state $|{\Phi_{0}}\rangle$ is a common zero-energy ground state. Hence, the 2D version of the fully polarized state $|{\Phi_{0}}\rangle\coloneqq\bigotimes_{x,y=1}^{L}|0\rangle_{(x,y)}$ is a simultaneous ground state of all terms in $\hat{H}$, implying that $\hat{H}$ is FF.
@@
To construct a variational state for low-energy excitations, let us define the plane-wave state of a single spin flip by $|\Psi_{\bm{k}}\rangle\coloneqq\frac{1}{L}\sum_{x,y=1}^{L}e^{i\bm{k}\cdot\bm{r}}\hat{s}_{(x,y)}^{-}|{\Phi_{0}}\rangle$. The variational energy is given by
$E(\bm{k})$ $\coloneqq\langle\Psi_{\bm{k}}|\hat{H}|\Psi_{\bm{k}}\rangle=E(k_{x})+E(k_{y}),$ (17)